# Donghyeon Lee1, Minbyul Jeong2∗, Jaewoo Kang1,3

### v1

@@ A.2.1 Biomedical Domain
A.2.2 Legal Domain
@@
A.2.3 Commonsense and mathematics
@@ A.3 Inference Setting
A.4 Tendency of Legal Domain and Time-Invariant Knowledge
Figure 6: Performance analysis of legal domain. Among time variant domains, legal domain shows the most stable results of static, besides, there still exists fluctuation and steep cutoff in dynamic dataset. When it comes to each templates, generation shows a less variation among models, while MCQA setting shows cut-off in many models between 2015 and 2016. It is also observed that in this domain, model’s capability of task is still influential such as the result of gemma-2-9b-it or SOLAR-10.7B-instruct-v1.0, but the parametric knowledge seems more important as small model like Phi-3.5-mini-instruct cannot overcome the gap.
Figure 7: Performance analysis of common-sense and mathematics domains. All model shows clearly the domain specific characteristics, which is invariant knowledge even it comes with temporal attributes.
Figure 6 shows the result of legal domain. Among time variant domains, legal domain shows the most stable results of static, also minimal decline in dynamic dataset, especially generation setting. This indicates the domain specificity, which has less frequent yearly changes and change continues across many time stamps. In the template wise viewpoint, however, MCQA templates exhibit cutoffs between 2015 and 2016, suggesting selection tasks are more sensitive in this domain’s temporal boundaries. Also, model’s capability for task is still influential in legal domain like the result of gemma-2-9b-it and SOLAR-10.7B-instruct-v1.0 shows, but the parametric knowledge seems more important here as small model cannot overcome the gap by technical skills, like Phi-3.5-mini-instruct.
For commonsense and mathematics in Figure 7, arbitrary years based on the biomedical domain were used, from 2020 to 2024. The left side of result shows the tendency of generation templates, and the right side is the tendency of MCQA templates, in both domains. Results show minimal variation, aligning with the stable nature of these knowledge types. This consistency confirms that time-invariant knowledge is well-preserved across models. For template wise comparison, generation cases show a way little gap between models, while MCQA tasks show the different between models, which is aligned with the findings from other time variant domains: the ability of each model’s specialized task affects its knowledge recall ability.
About the overall performance quality, the result of time invariant shows lower performance as the model show generate appropriate object in each knowledge, even though the time invariant knowledge’s coverage is wider than other domains. This tendency is alleviated by using MCQA templates, which ends of the rationale for using multiple templates.
A.5 Algorithm of ChroKnowPrompt
@@ 35   Update knowledge categorization to Chrono-Correct
A.6 Evaluation of ChroKnowPrompt in Legal domain
@@
Table 4 shows the evaluation of ChroKnowPrompt in legal domain. While our approach demonstrates significant improvements in certain domains, it shows limited or negligible gains in the legal domain. Overall score in dynamic dataset is mimic, as the highest gain is 1.9 increasing. However, the increase in static dataset is still impressive as the highest increase is comparable level as much as biomedical domain’s results (more than 10 in the proprietary model). Another finding is that although the increase in Table 3’s result in general domain is not higher than the static figures in the legal domain, the variation in figures between models is significantly larger in the legal domain. As the format of legal dataset is the unstructured format with long context, this would be one factor of low edit quality.
@@
Table 4: Result of ChroKnowPrompt in Legal domain. The order of open-source LLMs follows the same sequence as in Table 3, starting with the latest model and progressing to the most outdated one. The numeric score represents the level of Known in chronological categorization, and the increase in parentheses reflects the ratio of Chrono-correct answers, considering total span in the left side and previous span in the right side.
@@
Temporal Increase 0.4 4.9 0.3 3.5
Domain Increase 2.7 1.9

### v2

@@ A.3.4 Source and Approach of Biomedical Domain
A.3.5 Source and Approach of Legal Domain
@@
A.3.6 Source and Approach of Commonsense and mathematics
@@ A.4 Inference Setting
A.5 Details of ChroKnowledge in Legal and Time-Invariant Domain
Figure 9: Performance analysis of common-sense and mathematics domains. Three line plots represent each template’s results: Generation, MCQA and TF. All model shows clearly the domain specific characteristics, which is invariant knowledge even it comes with temporal attributes. Overall results are lower in generation templates, as it is challenging for models to correctly recall exactly one object in these domains (e.g., ’subject’: ’Parent’, ’relation’: ’Synonym’ has more objects later than ’Ancestor’)
Figure 4 shows the result of legal domain. Among time variant domains, legal domain shows the most stable results of static, also minimal decline in dynamic dataset. This indicates the domain specificity, which has less frequent yearly changes (almost cases has one change of object in total time frame) and change continues across many time stamps. Also, model’s capability for specific task setting is influential in legal domain like the result of MCQA and TF shows, where gap between generation and other templates are many times larger than in other domains.
For commonsense and mathematics in Figure 9, arbitrary years based on the biomedical domain were used, from 2020 to 2024. The left side of result shows the tendency of generation templates, and the middle side is the tendency of MCQA templates, and the last one for TF templates. Each results measure the percentage of Correct answer, represented as line plots. Results show minimal variation, aligning with the stable nature of these knowledge types. This consistency confirms that time-invariant knowledge is well-preserved across models. For template wise comparison, generation cases show a way little gap between models, while MCQA tasks show the different between models, which is aligned with the findings from other time variant domains: the ability of each model’s specialized task affects its knowledge recall ability.
About the overall performance quality, the result of time invariant shows lower performance as the models generate one object per each knowledge, while the time invariant knowledge’s coverage is wider than other domains. This tendency is alleviated by using MCQA and TF templates, which ends of the rationale for helpfulness of using multiple templates to check knowledge.
A.6 Total Time Frame Result of ChroKnowledge
Figure 10 represents the total results for all the range of time frames in general, and legal domain. Figure 11 to Figure 15 represents the total results of template-wise performance in ChroKnowledge. Each domain’s result is separated into two category of temporal state: Dynamic and Static. Every spider plots consist with three template: Generation, Multi-choice QA, and True/False. Each statistics refer the percentage of Correct answer, same as Figure 2.
A.7 Algorithm of ChroKnowPrompt
@@ 35   Update knowledge categorization to Chrono-Correct
A.8 Task Configurations of ChroKnowPrompt
We apply our method to both Incorrect and Partial Correct categories, as the latter may still lack definitive answers. The test set consists of 10% of the total dataset from each domain. Evaluation employs fuzzy matching with a temperature of 0 for strict assessment, classifying an answer as Chrono-correct only if the last candidate answer matches the object. As described in Section 6.2, the system prompt for each case (generation or verification&refinement) works as follows Table 4:
Generation Case
[System]
Answer ’Candidate A. [Object]’ based on the timestamp. Output only the answer: ’A. [Object]’.
Verify & Refine Case
[System]
Answer ’Candidate A. [Object]’ based on the timestamp. If it is correct, repeat the same [Object]. If it is wrong, generate a new [Object]. Output only the answer: ’A. [Object]’.
Table 4: System prompts for ChroKnowPrompt.
Table 5: Result of ChroKnowPrompt for both object changed and unchanged cases. The order of open-sources LLM is sorted by release date, starting from the latest model to the most outdated model. The numeric score represents the level of Known increase in chronological categorization, and the increase is due to the transition of previously confusing Partial correct responses to Chrono-correct. The parenthesis score is the total percentage in dynamic or static dataset, including both changed and unchanged cases. Showing almost 10% to 30% performance of object unchanged cases, the results gives observations that only prompting method has limitations in editing diversely changing temporal knowledge.
Models general biomedical legal
dynamic static dynamic static dynamic static
Object changed unchanged unchanged changed unchanged unchanged changed unchanged unchanged
Proprietary Large Language Models
GPT4o-mini +0.7 (28.7) +7.0 (28.7) +4.7 (33.2) +1.1 (51.9) +21.9 (51.9) +27.8 (51.6) +0.0 (3.2) +1.9 (3.2) +14.1 (51.9)
Gemini-1.5-flash +0.6 (15.6) +5.9 (15.6) +4.5 (22.1) +0.8 (49.0) +15.0 (49.0) +16.0 (48.8) +0.0 (1.3) +1.0 (1.3) +14.1 (16.3)
Open-Source Large Language Models
Phi3.5 Mini +0.3 (17.3) +1.8 (17.3) +2.5 (25.5) +2.1 (45.4) +16.7 (45.4) +20.3 (41.3) +0.0 (0.6) +0.3 (0.6) +4.5 (14.2)
LLaMA3.1 +0.2 (20.6) +2.8 (20.6) +1.7 (27.1) +1.4 (36.9) +7.8 (36.9) +7.9 (33.6) +0.0 (0.3) +0.0 (0.3) +1.3 (13.8)
Gemma2 +1.0 (19.6) +3.0 (19.6) +2.3 (26.7) +0.6 (32.5) +5.6 (32.5) +9.0 (31.7) +0.0 (2.9) +0.6 (2.9) +2.6 (44.6)
Mistral v0.3 +0.4 (18.6) +1.5 (18.6) +1.6 (26.9) +0.4 (26.6) +3.8 (26.6) +5.6 (24.3) +0.0 (1.3) +0.6 (1.3) +7.0 (21.1)
LLaMA3 +0.4 (20.9) +2.3 (20.9) +1.7 (28.0) +0.3 (31.4) +5.5 (31.4) +3.8 (25.7) +0.0 (1.0) +0.3 (1.0) +0.6 (18.9)
Gemma +0.2 (18.9) +0.8 (18.9) +1.5 (25.9) +0.3 (18.3) +5.7 (18.3) +5.3 (12.6) +0.0 (0.3) +0.0 (0.3) +0.0 (8.70)
SOLAR +0.1 (16.5) +0.7 (16.5) +0.9 (24.9) +0.3 (26.5) +3.8 (26.5) +4.5 (20.3) +0.0 (0.6) +0.0 (0.6) +1.3 (26.8)
LLaMA2 +0.3 (18.1) +4.9 (18.1) +5.0 (26.6) +2.0 (44.3) +23.2 (44.3) +26.3 (37.2) +0.0 (0.3) +0.0 (0.3) +12.8 (21.8)
Object Increase 0.4 2.9 0.9 11.8 0.0 3.2
Temporal Increase 1.7 2.6 5.9 12.7 0.2 5.8
Domain Increase 2.0 8.2 2.1
A.9 Details in span-wise results of ChroKnowPrompt
@@
Table 6 and 7 present the evaluation of ChroKnowPrompt in span-wise comparisons. While our approach demonstrates significant improvements in certain domains, it shows limited or negligible gains in the legal domain. Overall scores in the dynamic dataset remain modest, with the highest gain being only 1.9. However, the static dataset yields more impressive results, with the highest increase exceeding 10% in proprietary models, a level comparable to the biomedical domain’s results. Another finding is that although the increase in Table 6’s result in general domain is not higher than the static figures in the legal domain, the variation in figures between models is significantly larger in the legal domain. As the format of legal dataset is the unstructured format with long context, this would be one factor of low edit quality.
@@
Table 6: Result of ChroKnowPrompt in span-wise comparison for general and biomedical domain. The order of open-sources LLM is sorted by release date, starting from the latest model to the most outdated model. The numeric score is the level of Known in chronological categorization and the increase in parentheses is from the ratio of Chrono-correct which was confusing Partial correct before. Each result presents both in total span and previous span.
Models general biomedical Model Increase
total span previous span total span previous span total previous
Proprietary Large Language Models
Gemini-1.5-flash 15.6 (+6.5) 22.1 (+4.5) 15.3 (+6.1) 21.7 (+4.1) 49.0 (+15.8) 48.8 (+16.0) 48.0 (+14.9) 51.7 (+18.8) 10.7 11.0
Open-Source Large Language Models
Open-Source Chat Models
Mpt 18.3 (+4.8) 25.6 (+4.8) 17.0 (+3.5) 22.8 (+2.1) 43.3 (+22.9) 45.3 (+30.3) 30.8 (+10.4) 26.6 (+11.6) 15.7 6.9
Pythia 13.8 (+0.0) 20.8 (+0.1) 13.8 (+0.0) 20.7 (+0.0) 13.1 (+0.0) 10.2 (+0.1) 13.1 (+0.0) 10.2 (+0.1) 0.1 0.0
Nemotron3 11.2 (+1.5) 18.3 (+1.8) 10.1 (+0.5) 16.7 (+0.1) 22.1 (+9.0) 19.4 (+8.3) 17.9 (+4.8) 15.4 (+4.4) 5.2 2.5
Temporal Increase 3.2 2.5 2.3 1.4 11.5 12.7 6.6 6.8
Domain Increase 2.9 1.8 12.1 6.8
Table 7: Result of ChroKnowPrompt in span-wise comparison for legal domain. The order of open-source LLMs follows the same sequence as in Table 6, starting with the latest model and progressing to the most outdated one. The numeric score represents the level of Known in chronological categorization, and the increase in parentheses reflects the ratio of Chrono-correct answers, considering total span in the left side and previous span in the right side.
@@
Open-Source Chat Models
Mpt 1.0 (+0.6) 8.4 (+5.1) 0.6 (+0.3) 4.5 (+1.3) 2.9 0.8
Pythia 0.3 (+0.0) 3.2 (+0.0) 0.3 (+0.0) 3.2 (+0.0) 0.0 0.0
Nemotron3 0.3 (+0.0) 5.1 (+1.0) 0.3 (+0.0) 4.8 (+0.6) 0.5 0.3
Temporal Increase 0.4 5.0 0.3 3.8
Domain Increase 2.7 2.1
Figure 10: Total result of performance heatmap in general and legal domain
Figure 11: Total result of template-wise performance in general domain, dynamic dataset
Figure 12: Total result of template-wise performance in general domain, static dataset
Figure 13: Total result of template-wise performance in biomedical domain, dynamic and static dataset
Figure 14: Total result of template-wise performance in legal domain, dynamic dataset
Figure 15: Total result of template-wise performance in legal domain, static dataset
@@ A. United States Secretary of the Navy
Template: TF
Q. In 2020, Pedro Braillard Poccard, position held, member of the Argentine Chamber of Senators
A. true
Q. In 2020, Ana Santos Aramburo, position held, Director of the Prado Museum
A. false
Q. In 2020, James E. McPherson, position held, United States Secretary of Defense
A. false
Q. In 2020, Jesús Ávila de Grado, position held, chief scientific officer
A. true
Q. In 2020, Donald Tusk, position held, Prime Minister of Poland
: (generate from here …)
Table 9: Example of TF templates. Domain: general, Subject: Donald Tusk, Relation: position held (P39), year to target: 2020.